OpenStudy (anonymous):

Identify the GCF of 16x^4y^3 + 8x^3y^2 - 20x^2y.

OpenStudy (anonymous):

GCF is Greatest Common Factor

OpenStudy (apoorvk):

Man man, what do you find common in the three terms in the expression - \[16x^4y^3 , 8x^3y^2 - 20 x^2y\]

OpenStudy (anonymous):

x and y after every number

OpenStudy (apoorvk):

well yeah but umm... Okay let's do it this way: \[16x^4y^3 = 2 \times 2 \times 2 \times 2 \times x \times x \times x\times x\times y \times y \times y\]\[8x^3y^2 = 2 \times 2 \times 2 \times x \times x \times x \times y \times y \]\[20x^2y = 2 \times 2 \times 5 \times x \times x \times y\] i listed the factors of the three terms - what's common to all three of 'em?

OpenStudy (anonymous):

I dont see it